Cash Payback Period, Net Present Value Method, and Analysis

Elite Apparel Inc. is considering two investment projects. The estimated net cash flows from each project are as follows:

Year Plant Expansion Retail Store Expansion
1 $145,000 $121,000
2 119,000 143,000
3 102,000 98,000
4 93,000 68,000
5 29,000 58,000
Total $488,000 $488,000

Each project requires an investment of $264,000. A rate of 10% has been selected for the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the cash payback period for each project.

Cash Payback Period
Plant Expansion 2 years
Retail Store Expansion 2 years

1b. Compute the net present value. Use the present value of $1 table above. If required, round to the nearest dollar.

Plant Expansion Retail Store Expansion
Present value of net cash flow total $ $
Less amount to be invested $ $
Net present value $ $
Answer #1

Answer- 1-b)- Net present value - Plant Expansion= $124229

Retail store expansion = $120167

Explanation-

Calculation of Investment's Net Present Value
Project Plant Expansion
Net Cash Flows $ (a) Present Value of 1 at 10% (b) Present Value of cash flows (c=a*b) $
Year 1 145000 0.909 131805
Year 2 119000 0.826 98294
Year 3 102000 0.751 76602
Year 4 93000 0.683 63519
Year 5 29000 0.621 18009
Totals
Total present value of cash inflow (a) 388229
Total cash outflow (b) -264000 1 -264000
Net Present Value $ (c=a-b) 124229
Calculation of Investment's Net Present Value
Project Retail Store Expansion
Net Cash Flows $ (a) Present Value of 1 at 10% (b) Present Value of cash flows (c=a*b) $
Year 1 121000 0.909 109989
Year 2 143000 0.826 118118
Year 3 98000 0.751 73598
Year 4 68000 0.683 46444
Year 5 58000 0.621 36018
Totals
Total present value of cash inflow (a) 384167
Total cash outflow (b) -264000 1 -264000
Net Present Value $ (c=a-b) 120167